# Bringing memory to Boolean networks: a unifying framework

### v1

@@ 3.2 Refinement of the hierarchy by reachability
The hierarchy by reachability gives $\mathtt{C}^{*}\subseteq\mathtt{M}^{*}$ and $\mathtt{H}^{*}\subseteq\mathtt{M}^{*}$, or in other words a reachability pair that is either history-based or cuttable is most permissive. We now prove that there are most permissive reachability pairs that are neither history-based nor cuttable, i.e. $\mathtt{M}^{*}\neq(\mathtt{H}^{*}\cup\mathtt{C}^{*})$. Let $\mathsf{C}\in\mathrm{F}(3)$ from Example 3.9 and $\mathsf{H}$ from Example 3.8, and let $f\in\mathrm{F}(6)$ be defined by
@@ 3.3 Equivalent update modes
We now classify the commensurate and min-trapspace-equivalent update modes from our list. Since $\mathtt{T}^{*}=\mathtt{S}^{*}$, we omit $\mathtt{S}$.
@@
We first prove that $\mathtt{M}$ is min-trapspace-equivalent to $\mathtt{T}$.
@@
For any configuration $x$ and any min-trapspace configuration $y\in T_{f}(x)$, we have $x\to_{\mathtt{M}}^{*}y$.
@@
We first prove that $x\to_{\mathtt{M}}^{*}u:=T_{f}(x)-x$. Let $z\in T_{f}(x)$ be the furthest configuration reachable from $x$ in $\mathtt{M}$, and let $x=x^{0}\to_{\mathtt{M}}\dots\to_{\mathtt{M}}x^{l-1}=z$. For the sake of contradiction, assume $z\neq u$. Since $[x,z]$ is a strict subcube of $T_{f}(x)$, it is not a trapspace, hence there exists $s^{l}\in[x,z]\subseteq[x^{0},\dots,x^{l-1}]$ such that $f(s^{l})\notin[x,z]$. In particular, there exists a coordinate $i^{l}\in[n]$ such that $x_{i^{l}}=z_{i^{l}}\neq f_{i^{l}}(s^{l})$. Letting $w^{l}=(i^{l},s^{l},t^{l}=z)$, we obtain $x^{l}$ that is reachable from $x$ and yet is further away from $x$ than $z$ is, which is the desired contradiction.
@@
For any configuration $x$ with $|\mathtt{T}^{*}(x)|=2^{d}$, then $|\mathtt{M}^{*}(x)|\geq L(d)=2^{\lfloor d/2\rfloor}+2^{\lceil d/2\rceil}-1$. Conversely, for any $L(d)\leq k\leq 2^{d}$, there exists $f\in\mathrm{F}(d)$ and $x\in\mathbb{B}^{d}$ such that $T_{f}(x)=\mathbb{B}^{d}$ while $|\mathtt{M}^{*}(x)|=k$.
@@
Without loss, let $d=n$ and $x=0\dots 0$. Let $C$ be the set of local functions $f_{i}$ that are constant (equal to $1$) in $T_{f}(x)=\mathbb{B}^{n}$, and let $c=|C|$ and $D=[n]\setminus C$. Then $x$ can always reach the following two sets of configurations in $\mathtt{M}$.
@@
We prove, by induction on $n$, that for any $L(n)\leq k\leq 2^{n}$ there exists $f\in\mathrm{F}(n)$ and $x\in\mathbb{B}^{n}$ such that $T_{f}(x)=\mathbb{B}^{n}$ and $|\mathtt{M}^{*}(x)|=k$. We also use its immediate consequence: for any $n\geq 1$ and any $1\leq r\leq 2^{n}$, there exists $f\in\mathrm{F}(n)$ and $x\in\mathbb{B}^{n}$ such that $|\mathtt{M}^{*}(x)|=r$. The claim is trivial for $n=1$ so let us assume it holds up to $n-1$. Let $c=\lfloor n/2\rfloor$ and define $q$ and $r$ as the quotient and remainder of the following long division:
@@
By induction hypothesis, let $g\in\mathrm{F}(n-c)$ and $z=0\dots 0\in\mathbb{B}^{n-c}$ such that $|\mathtt{M}^{*}(z)|=r$. Let $C=\{1,\dots,c\}$ and $D=\{c+1,\dots,n\}$ and let $f\in\mathrm{F}(n)$ be defined as
@@
$f_{D}(x)$ =$=\begin{cases}\neg x_{D}&\text{if }x_{C}\in Q\setminus\{q^{*}\}\\ g(x_{D})&\text{if }x_{C}=q^{*}\\ x_{D}&\text{if }x_{D}\notin Q.\end{cases}$
@@
We can now determine the configurations reachable from $x$. Firstly, if $\alpha\in\mathbb{B}^{c}\setminus Q$, then $f_{D}(\alpha,x_{D})$ reduces to the identity, hence $R_{\alpha}=\{(\alpha,z)\}$. Secondly, if $\alpha=q^{*}$, then $f_{C}(\alpha,x_{D})$ reduces to $g$, hence $R_{q^{*}}=\{(q^{*},b):b\in\mathtt{M}^{*}(z)\}$. Thirdly, if $\alpha\in Q\setminus\{q^{*}\}$, then $f_{D}(\alpha,x_{D})$ reduces to the negation, hence $R_{\alpha}=\{(\alpha,a):a\in\mathbb{B}^{n-c}\}$. Therefore, $x$ reaches the following set of configurations:
$\mathtt{M}^{*}(x)$ =$=\{(\alpha,z):\alpha\in\mathbb{B}^{c}\setminus Q\}\cup\{(q^{*},b):b\in\mathtt{M}^{*}(z)\}\cup\{(\alpha,a):\alpha\in Q\setminus\{q^{*}\},a\in\mathbb{B}^{n-c}\},$
@@
We now exhibit some min-trapspace configurations that are reachable in some update mode $\mu$ but not in another mode $\nu$. Those examples also show that $\mu$ and $\nu$ are not commensurate. The three examples below follow a similar structure; as such, we only provide a detailed explanation for the first example.
@@
Let $\mathsf{H}\in\mathrm{F}(3)$ be the network from Example 2.2 and let $s=101$. It is easily seen that there is no $\mathtt{C}$-trajectory such that $s^{a}=s$. Then let $\hat{\mathsf{H}}\in\mathrm{F}(n)$ be defined as
$\hat{\mathsf{H}}(x)=\begin{cases}(\neg x_{123},\neg x_{4..n-1},1)&\text{if }x_{n}=1\\ (\mathsf{H}(x_{123}),0_{4..n-1},0)&\text{if }x_{n}=0\text{ and }x_{4..n-1}=0\text{ and }x_{123}\neq s\\ (\mathsf{H}(s),0_{4..n-1},1)&\text{if }x_{n}=0\text{ and }x_{4..n-1}=0\text{ and }x_{123}=s\\ x&\text{otherwise}.\end{cases}$
Note that there is only one transition from the hyperplane $\{x\in\mathbb{B}^{n}:x_{n}=0\}$ to its parallel $\{x\in\mathbb{B}^{n}:x_{n}=1\}$. Then $\hat{x}=0\dots 0$ can reach the whole of $\{x\in\mathbb{B}^{n}:x_{n}=0\}$ in $\mathtt{H}$ but it can only reach as most the seven configurations $\{(y_{123},0_{4..n-1},0):y\neq s\}$ in $\mathtt{C}$.
Example 3.15 (Min-trapspace configurations reachable in cuttable but not in history-based or interval).
Let $\mathsf{C}\in\mathrm{F}(3)$ be the network from Example 2.7 and $s=010$, then again no history-based or interval trajectory can have $s^{a}=s$. Then let $\hat{\mathsf{C}}\in\mathrm{F}(n)$ be defined as
$\hat{\mathsf{C}}(x)=\begin{cases}(\neg x_{123},\neg x_{4..n-1},1)&\text{if }x_{n}=1\\ (\mathsf{C}(x_{123}),0_{4..n-1},0)&\text{if }x_{n}=0\text{ and }x_{4..n-1}=0\text{ and }x_{123}\neq s\\ (\mathsf{C}(s),0_{4..n-1},1)&\text{if }x_{n}=0\text{ and }x_{4..n-1}=0\text{ and }x_{123}=s\\ x&\text{otherwise}.\end{cases}$
The two previous examples also prove that $\mathtt{M}$ is not min-trapspace equivalent to either $\mathtt{C}$ or $\mathtt{H}$.
@@
Let $\mathsf{I}\in\mathrm{F}(2)$ be the network from Example 3.10. Let $s=01$ and $\hat{\mathsf{I}}\in\mathrm{F}(n)$ be defined as
$\hat{\mathsf{I}}(x)=\begin{cases}(\neg x_{12},\neg x_{3..n-1},1)&\text{if }x_{n}=1\\ (\mathsf{I}(x_{12}),0_{3..n-1},0)&\text{if }x_{n}=0\text{ and }x_{3..n-1}=0\text{ and }x_{12}\neq s\\ (\mathsf{I}(s),0_{3..n-1},1)&\text{if }x_{n}=0\text{ and }x_{3..n-1}=0\text{ and }x_{12}=s\\ x&\text{otherwise}.\end{cases}$
4 Principal trapspaces
4.1 Principal trapspaces and trapping closure
The general asynchronous graph of a BN $f\in\mathrm{F}(n)$ is the graph $\mathtt{GA}(f)=(V,E)$ where $V=\mathbb{B}^{n}$ and
$E=\{(x,f^{(S)}(x)):x\in\mathbb{B}^{n},S\subseteq[n]\}.$
It is clear that the mapping $\mathtt{GA}:f\mapsto\mathtt{GA}(f)$ is injective. We first characterise the general asynchronous graphs.
@@ 4.1 Principal trapspaces and trapping closure
Let $\Gamma$ be a graph on $\mathbb{B}^{n}$. Then $\Gamma=\mathtt{GA}(f)$ for some $f\in\mathrm{F}(n)$ if and only if $\Gamma$ is reflexive and all out-neighbourhoods are subcubes.

### v2

@@ 3.2 Refinement of the hierarchy by reachability
The hierarchy by reachability gives $\mathtt{C}^{*}\subseteq\mathtt{M}^{*}$ and $\mathtt{H}^{*}\subseteq\mathtt{M}^{*}$, or in other words a reachability pair that is either history-based or cuttable is most permissive. We now prove that there are most permissive reachability pairs that are neither history-based nor cuttable, i.e. $\mathtt{M}^{*}\neq(\mathtt{H}^{*}\cup\mathtt{C}^{*})$. Let $\mathsf{C}\in\mathrm{F}(3)$ from the proof of (i) and $\mathsf{H}$ from the proof of (h), and let $f\in\mathrm{F}(6)$ be defined by
@@ 3.3 Equivalent update modes
We now classify the commensurate and min-trapspace-equivalent update modes from our list, and show these two notions are equivalent. Since $\mathtt{T}^{*}=\mathtt{S}^{*}$, we omit $\mathtt{S}$.
@@
We first prove that $\mathtt{M}$ and $\mathtt{T}$ are min-trapspace-equivalent (Lemma 3.9) and commensurate (Proposition 3.10).
Then, using examples, we show that: (1) that the most permissive/history-based update mode is neither min-trap-equivalent nor commensurate to the cuttable/interval update mode. (Example 3.12); (2) the most permissive/cuttable update mode is neither min-trap-equivalent nor commensurate to the history-based/interval update mode (Example 3.13); (3) the interval and asynchronous update modes are neither min-trapspace-equivalent nor commensurate (Example 3.14).
@@
For any configuration $x$ and any min-trapspace configuration $y\in T_{f}(x)$ (i.e., $x\to^{*}_{\mathtt{T}}y$), we have $x\to_{\mathtt{M}}^{*}y$.
@@
We first prove that $x\to_{\mathtt{M}}^{*}u:=T_{f}(x)-x$. Let $z\in T_{f}(x)$ be the furthest configuration reachable from $x$ in $\mathtt{M}$, and let $x=x^{0}\to_{\mathtt{M}}\dots\to_{\mathtt{M}}x^{l-1}=z$. For the sake of contradiction, assume $z\neq u$. Since $[x,z]$ is a strict subcube of $T_{f}(x)$, it is not a trapspace, hence there exists $s^{l}\in[x,z]\subseteq[x^{0},\dots,x^{l-1}]$ such that $f(s^{l})\notin[x,z]$. In particular, there exists a component $i^{l}\in[n]$ such that $x_{i^{l}}=z_{i^{l}}\neq f_{i^{l}}(s^{l})$. Letting $w^{l}=(i^{l},s^{l},t^{l}=z)$, we obtain $x^{l}$ that is reachable from $x$ and yet is further away from $x$ than $z$ is, which is the desired contradiction.
@@
For any configuration $x$ with $|\mathtt{T}^{*}(x)|=2^{d}$, we have $|\mathtt{M}^{*}(x)|\geq L(d)=2^{\lfloor d/2\rfloor}+2^{\lceil d/2\rceil}-1$. Conversely, for any $L(d)\leq k\leq 2^{d}$, there exists $f\in\mathrm{F}(d)$ and $x\in\mathbb{B}^{d}$ such that $T_{f}(x)=\mathbb{B}^{d}$ while $|\mathtt{M}^{*}(x)|=k$.
@@
Without loss, let $d=n$ and $x=0^{n}$. Let $C$ be the set of local functions $f_{i}$ that are constant (equal to $1$) in $T_{f}(x)=\mathbb{B}^{n}$, and let $c=|C|$ and $D=[n]\setminus C$. Then $x$ can always reach the following two sets of configurations in $\mathtt{M}$.
@@
We prove, by induction on $n$, that for any $L(n)\leq k\leq 2^{n}$ there exists $f\in\mathrm{F}(n)$ and $x\in\mathbb{B}^{n}$ such that $T_{f}(x)=\mathbb{B}^{n}$ and $|\mathtt{M}^{*}(x)|=k$. We also use its immediate consequence: for any $n\geq 1$ and any $1\leq r\leq 2^{n}$, there exists $f\in\mathrm{F}(n)$ and $x\in\mathbb{B}^{n}$ such that $|\mathtt{M}^{*}(x)|=r$. The claim is trivial for $n=1$ so let us assume it holds up to n1n-1. Let $c=\lfloor n/2\rfloor$ and define $q$ and $r$ as the quotient and remainder of the following long division:
@@
By induction hypothesis, let $g\in\mathrm{F}(n-c)$ and $z=0^{n-c}\in\mathbb{B}^{n-c}$ such that $|\mathtt{M}^{*}(z)|=r$. Let $C=\{1,\dots,c\}$ and $D=\{c+1,\dots,n\}$ and let $f\in\mathrm{F}(n)$ be defined as
@@
$f_{D}(x)$ $=\begin{cases}\neg x_{D}&\text{if }x_{C}\in Q\setminus\{q^{*}\}\\ g(x_{D})&\text{if }x_{C}=q^{*}\\ x_{D}&\text{if }x_{C}\notin Q.\end{cases}$
@@
We can now determine the configurations reachable from $x$. Firstly, if $\alpha\in\mathbb{B}^{c}\setminus Q$, then $f_{D}(\alpha,x_{D})$ reduces to the identity, hence $R_{\alpha}=\{(\alpha,z)\}$. Secondly, if $\alpha=q^{*}$, then $f_{D}(\alpha,x_{D})$ reduces to $g$, hence $R_{q^{*}}=\{(q^{*},b):b\in\mathtt{M}^{*}(z)\}$. Thirdly, if $\alpha\in Q\setminus\{q^{*}\}$, then $f_{D}(\alpha,x_{D})$ reduces to the negation, hence $R_{\alpha}=\{(\alpha,a):a\in\mathbb{B}^{n-c}\}$. Therefore, $x$ reaches the following set of configurations:
$\mathtt{M}^{*}(x)$ $=\{(\alpha,z):\alpha\in\mathbb{B}^{c}\setminus Q\}\cup\{(q^{*},b):b\in\mathtt{M}^{*}(z)\}\cup\{(\alpha,a):\alpha\in Q\setminus\{q^{*}\},a\in\mathbb{B}^{n-c}\},$
@@
Corollary 3.11.
For any configuration $x$, we have
$|\mathtt{M}^{*}(x)|\leq|\mathtt{T}^{*}(x)|,\qquad|\mathtt{T}^{*}(x)|\leq(|\mathtt{M}^{*}(x)|)^{2}.$
Therefore, the most permissive and trapping update modes are commensurate.
We now exhibit three examples of min-trapspace configurations that are reachable in some update mode $\mu$ but not in another mode $\nu$; these three examples also show that $\mu$ and $\nu$ are not commensurate. Since the three examples follow a similar structure, we only provide a detailed explanation for the first example. We use the following shorthand notation: for all integers $i\leq j$, let $i..j=\{i,i+1,\dots,j\}$.
@@
Let $\mathsf{H}\in\mathrm{F}(3)$ be the network from Example 2.2 and let $s=101$. It is easily seen that there is no $\mathtt{C}$-trajectory starting at $x^{0}=000$ such that $s^{a}=s$. Then, for any $n\geq 5$, let $\hat{\mathsf{H}}\in\mathrm{F}(n)$ be defined as
[…]
Note that there is only one transition from the hyperplane $I=\{x\in\mathbb{B}^{n}:x_{n}=0\}$ to its parallel $J=\{x\in\mathbb{B}^{n}:x_{n}=1\}$. We note that $J$ is a minimal trapspace, as $T_{f}(y)=J$ for all $y\in J$. Then the configuration $\hat{x}=0^{n}$ can reach the whole of $J$ in $\mathtt{H}$ but it can only reach at most the seven configurations $\{(y_{123},0_{4..n-1},0):y\neq s\}$ in $\mathtt{C}$. Therefore, the history-based and cuttable update modes are not min-trapspace-equivalent.
This example also shows that the history-based and cuttable update modes are not commensurate. Indeed, we have
$|\mathtt{H}^{*}(\hat{x})|\geq 2^{n-1},\qquad|\mathtt{C}^{*}(\hat{x})|\leq 7.$
Thus, there is no function $\psi:\mathbb{N}\to\mathbb{N}$ such that $|\mathtt{H}^{*}(x)|\leq\psi(|\mathtt{C}^{*}(x)|)$ for all configurations $x$, as $\psi(7)$ would be unbounded.
[…]
Example 3.13 (Min-trapspace configurations reachable in cuttable but not in history-based).
Let $\mathsf{C}\in\mathrm{F}(3)$ be the network from Example 2.9 and $s=010$, then again no history-based or interval trajectory starting at $x^{0}=000$ can have $s^{a}=s$. Then, for any $n\geq 5$, let $\hat{\mathsf{C}}\in\mathrm{F}(n)$ be defined as
[…]
This example is a “mirror image” to Example 3.12. As such, this example shows that the most permissive/cuttable update mode is neither min-trap-equivalent nor commensurate to the history-based/interval update mode.
@@
Let $\mathsf{I}\in\mathrm{F}(3)$ be the network from Example 2.7 and $s=011$, then there again no asynchronous trajectory starting at $x^{0}=000$ can have $s^{a}=s$. Then, for any $n\geq 5$, let $\hat{\mathsf{I}}\in\mathrm{F}(n)$ be defined as
[…]
Similarly to the previous examples, this example shows that the interval update mode is neither min-trapspace-equivalent nor commensurate to the asynchronous update mode.
4 Consequences of using memory in update modes
In this subsection, we would like to highlight one important common point and three major differences between the asynchronous update mode $\mathtt{A}$ and any other update mode with memory $\mu\in\{\mathtt{H},\mathtt{T},\mathtt{M},\mathtt{S},\mathtt{I},\mathtt{C}\}$ discussed in this paper. The key observation is that the memory of the trajectory offers a context upon which more transitions are allowed. Therefore, when we arrive at a particular configuration $x^{a-1}$, one may usually perform more transitions than what would be allowed if the trajectory were starting from there, as given by the asynchronous graph. This yields the following four features of update modes that use memory. In the sequel, we assume $\mu\in\{\mathtt{H},\mathtt{T},\mathtt{M},\mathtt{S},\mathtt{I},\mathtt{C}\}$ and we fix the Boolean network $f\in\mathrm{F}(n)$.
  1. 1.
    Trajectories closed by concatenation. Note that in a general update mode $\nu$, as defined at the start of Section 2.1, we may have two consecutive $\nu$-trajectories $x^{0}\to_{\nu}\dots\to_{\nu}x^{l}$ and $x^{l}=y^{0}\to_{\nu}\dots\to_{\nu}y^{m}$ such that their concatenation $x^{0}\to\dots\to x^{l}=y^{0}\to\dots\to y^{m}$ is no longer a $\nu$-trajectory. This happens for instance if the update mode $\nu$ constrains the source $s^{a}$ and the target $t^{a}$ to be equal to the starting configuration of the trajectory. However, for any $\mu\in\{\mathtt{A},\mathtt{H},\mathtt{T},\mathtt{M},\mathtt{S},\mathtt{I},\mathtt{C}\}$, trajectories are closed under concatenation: if $x^{0}\to_{\mu}\dots\to_{\mu}x^{l}$ and $x^{l}=y^{0}\to_{\mu}\dots\to_{\mu}y^{m}$, then $x^{0}\to_{\mu}\dots\to_{\mu}x^{l}=y^{0}\to_{\mu}\dots\to_{\mu}y^{m}$. Intuitively, if the transitions $y^{a-1}\to y^{a}$ are allowed with only the memory from $x^{l}=y^{0}$, then they will still be allowed with an enhanced memory that goes back to $x^{0}$.
  2. 2.
    Trajectories not closed by suffix. Say a $\mu$-trajectory $(x^{0},\dots,x^{l})$ for some update mode $\mu$ is suffix-closed if for all $1\leq a\leq l$, $(x^{a},\dots,x^{l})$ is also a $\mu$-trajectory. Asynchronous trajectories are suffix-closed, i.e. if $x^{0}\to_{\mathtt{A}}x^{1}\to_{\mathtt{A}}\dots\to_{\mathtt{A}}x^{l}$, then for all $1\leq a\leq l$, $x^{a}\to_{\mathtt{A}}\dots\to_{\mathtt{A}}x^{l}$. This is no longer the case for any update mode $\mu$ with memory. In fact, any $\mu$-trajectory that is not an asynchronous trajectory (such a trajectory exists due to the hierarchy by trajectory in Figure 2) is not suffix-closed, as it takes advantage of memory to use a transition that is not allowed in the asynchronous case. For instance, the $\mathtt{H}$-trajectory in Example 2.2 is not suffix-closed: the suffix $(111,101)$ is not an $\mathtt{H}$-trajectory since $111$ is a fixed point.
  3. 3.
    No transition graph. This item is closely related to the previous one. The asynchronous update mode could be simply represented by a transition graph, namely the asynchronous graph $\mathtt{A}(f)$: $x^{0}\to_{\mathtt{A}}x^{1}\to_{\mathtt{A}}\dots\to_{\mathtt{A}}x^{l}$ if and only if $(x^{a-1},x^{a})$ is an arc in $\mathtt{A}(f)$. The same cannot be said for any other update mode discussed in this paper. Indeed, for any update mode with memory $\mu$, we have $x^{0}\to_{\mu}x^{1}$ if and only if $(x^{0},x^{1})$ is an arc in $\mathtt{A}(f)$; as such, the only possible choice for a $\mu$-transition graph” (i.e. a transition graph according to $\mu$) would be $\mathtt{A}(f)$. However, as shown in the hierarchy by trajectory, there exists a $\mu$-trajectory that is not asynchronous (see Example 2.7 for instance); hence the $\mu$-trajectories cannot be represented by a transition graph.
  4. 4.
    No clear definition of limit configurations and attractors. In the asynchronous update mode, there is a clear definition of an $\mathtt{A}$-limit configuration of $f$: those are the configurations $y$ such that, if $y\to_{\mathtt{A}}^{*}x$, then $x\to_{\mathtt{A}}^{*}y$. Equivalently, they belong to the terminal strong components of the asynchronous graph. In particular, any fixed point is a limit configuration for the asynchronous update mode. An $\mathtt{A}$-attractor of $f$ is any set of configurations of the form $\mathtt{A}^{*}(y)$, where $y$ is a limit configuration; equivalently it is a terminal strong component of the asynchronous graph.
    One can generalise these definitions to any update mode $\mu$: say $y$ is a $\mu$-limit configuration of $f$ if $y\to_{\mu}^{*}x$ implies $x\to_{\mu}^{*}y$, and a $\mu$-attractor of $f$ is any set of configurations of the form $\mu^{*}(y)$, where $y$ is a $\mu$-limit configuration. These definitions are mathematically valid, and indeed for any fixed point $y$, $y$ is a $\mu$-limit configuration and $\{y\}$ is a $\mu$-attractor. Moreover, Lemma 3.9 yields the following characterisation of $\mu$-attractors for $\mu\in\{\mathtt{M},\mathtt{T},\mathtt{S}\}$.
@@ 4 Consequences of using memory in update modes
    Let $f\in\mathrm{F}(n)$ and $\mu\in\{\mathtt{M},\mathtt{T},\mathtt{S}\}$. Then the $\mu$-limit configurations of $f$ are the min-trapspace configurations and the $\mu$-attractors of $f$ are its minimal trapspaces.